OpenStudy (josephreak):

Can't solve this on my TI 89 Titanium. The correct answer is -1 Question: lim as x approaches 90 from the right: cos(x)/x-90 How can I solve this on my TI 89? It gives me undefined, it just plugs the values in I guess..?

myininaya (myininaya):

is that \[\lim_{x \rightarrow 90^+} \frac{\cos(x)}{x-90} ?\] is this in degrees?

OpenStudy (josephreak):

Yes, we've always been using degrees.

myininaya (myininaya):

does your calculator know you are working in degrees?

OpenStudy (josephreak):

Yes, I even tried out all the other modes out of frustration lol

myininaya (myininaya):

this question is weird if the answer really is suppose to be -1, there looks like there is a mixture of units going on .... Like if the question was \[\lim_{x \rightarrow \frac{\pi}{2}^+} \frac{\cos(x)}{x-\frac{\pi}{2}}\] this would make more sense

OpenStudy (josephreak):

Um, yes it's like that, but since we've always used it as 180/2 then I always assumed it's no difference if I assume it's 90

myininaya (myininaya):

do you mean the problem was like this and then you change the pi/2 to 90 degrees?

OpenStudy (josephreak):

yes

myininaya (myininaya):

that is probably why your calculator is spitting out undefined then

OpenStudy (josephreak):

hm, so how can I make it spit out the right answer?

myininaya (myininaya):

don't change pi/2 to 90

OpenStudy (josephreak):

and stay in the same mode? (Degrees)

myininaya (myininaya):

no you shouldn't be doing much with degrees in calculus it should be all about radians mostly

myininaya (myininaya):

the calculator should be in radians

OpenStudy (josephreak):

ok, I'll try and see. thanks for your help

OpenStudy (josephreak):

It worked!!! thank you!! So, I should now stay in Radians when there's always PI and change to degrees when specified otherwise, right?

myininaya (myininaya):

stay in radians i never seen a calculus class ever consider degrees you had x approaches pi/2 earlier you changed this to x approaches 90 deg you change the "units" for a number while not changing the units for x but I guess if you want to you could write \[\lim_{x \rightarrow \frac{\pi}{2}} \frac{\cos(x)}{x-\frac{\pi}{2}} =\lim_{x \rightarrow 90^o} \frac{\cos(\frac{x \pi}{180^o})}{\frac{ x \pi}{180^o}-\frac{\pi}{2}}\]

OpenStudy (josephreak):

ok thanks

myininaya (myininaya):

but yeah I wouldn't do that I wouldn't change radians to degrees ever in calculus